Parth (parthkohli):

If \[ 34 + 4 = 43\] \[ 34 \times 4 = ?\]

OpenStudy (rsadhvika):

base 5

Parth (parthkohli):

What are these numbers? base 5.. base 12... base 10?

Parth (parthkohli):

Base 5 is like 1,2,3,4,5,10

OpenStudy (ash2326):

We use 0-9 digits normally it's base 10 if a number system has digits from 0-3 only, it's base 4 0, 1, 2, 3, 10

OpenStudy (ash2326):

@ParthKohli base 5 will have digits from 0-4 so 0, 1, 2, 3, 4, 10

Parth (parthkohli):

34 x 4 = 34 + 34 + 34 + 34 base 5?

Parth (parthkohli):

How do we calculate \(34\) in base 5?

OpenStudy (ash2326):

yeah:) to easily do this, let's convert the number into base 10. Multiply it and then convert back to base 5

Parth (parthkohli):

How can we convert?

OpenStudy (ash2326):

If we have 25 it's actually \[25=5\times 10^0+2\times 10^1\] Isn't it?

Parth (parthkohli):

Yeah! :)

OpenStudy (ash2326):

34 in base 10 \[34=4*5^0+3*5^1\] what will you get?

Parth (parthkohli):

So \(34 = 4 \times 5^{0} \times 3 \times 5^{1}\)?

Parth (parthkohli):

Oops. I meant a + in between.

Parth (parthkohli):

4 + 15 = 19

OpenStudy (ash2326):

4 will be 4

OpenStudy (ash2326):

so \[19*4=??\]

Parth (parthkohli):

\(19 \times 4 = 76\)

Parth (parthkohli):

\(76 = 7 \times 10^1 + 6 \times 10^0\)

OpenStudy (ash2326):

now we have to convert into base 5 divide 76 by 5 15 ______________________ 5|76 75 first division result in quotient of 15 and remainder 1, so 1 is your first digit Now divide 15 by 5 and then tell me the remainder

Parth (parthkohli):

0 is the remainder.

Parth (parthkohli):

I get 301. Is that correct?

OpenStudy (ash2326):

Correct:D

Parth (parthkohli):

What? I just guessed it :P

OpenStudy (rsadhvika):

another way is you can do it directly in base 5 without converting

OpenStudy (ash2326):

you got remainder 1 then 0 and then 3 so it's 301

mathslover (mathslover):

\[\large{(10a+b)+b=10b+a}\] \[\large{10a+2b=10b+a}\] \[\large{9a=8b}\] \[\large{a=\frac{8b}{9}}\] \[\large{(10a+b)*b=10ab+b^2=10(\frac{8b}{9})+b^2=\frac{80b+9b^2}{9}}\] b = 4: \[\large{\frac{320+144}{9}=\frac{464}{9}}\]

mathslover (mathslover):

:d is that correct way @ash2326 ?

Parth (parthkohli):

How do you get 3? I just guessed it.

OpenStudy (asnaseer):

as @rsadhvika indicated, you can also multiply directly in base 5

Parth (parthkohli):

How to do so?

OpenStudy (ash2326):

when you divided by 15 you got 3 as quotient and remainder 0 next step is to divide 3 by 5 and then find the remainder

OpenStudy (asnaseer):

let me show you...

OpenStudy (ash2326):

\[3*5^2+0*5^1+1*5^0=76\]

OpenStudy (asnaseer):